# Annotator-Centric Active Learning for Subjective NLP Tasks

### v1

@@ 5 Results
5.1 Test sets results
See Figure 2 for the performance of the DICES, MFTC, and MHS, respectively. For MFTC, we initially focus on care, since it is the task with neither the most nor least amount of data. For MHS, we start with dehumanize, since it saw the most medium-level disagreement. The rest of the results can be observed in Appendix B.
Combining our results across datasets, we see that data characteristics influence whether ACAL can learn performant models efficiently. In particular, we see that for DICES and MHS, ACAL may learn models that perform well using less data (38% and 62% reduction at best, respectively). Conversely, for MFTC, there is little impact of using ACAL over PL (5.6% less data used). A similar pattern holds when comparing ACAL to AL, though AL seems to be a strong baseline for MHS, where random sample selection leads to more efficient data usage (60%). AL with uncertainty sampling is more efficient for MFTC (13%).
When we compare the performance metrics, we see that the distributions obtained through ACAL are consistently closer to the ground truth distribution in DICES, as measured by than PL and AL. However, this pattern is not visible for MFTC and MHS. In terms of majority-voted , ACAL again leads to better scores in both DICES and MHS. Since DICES and MHS are datasets with moderate disagreement, we may benefit from using ACAL in such scenarios. Further, if the dataset contains a large number of annotators per sample, annotator selection strategies are shown to pick a more informative set of annotators to learn from.
We highlight some further dataset-specific findings that shed light on the differences between the annotator selection strategies in ACAL. First, in DICES, we see that for three out of four annotator sampling strategies (, , ), the choice of data sampling strategy has no impact on the performance of the model due to the low number of samples to choose from. Furthermore, only performs worse in terms of , overrepresenting outlier annotators. This hints that selecting annotators based on the average embedding of the annotated content strongly emphasizes diverging label behavior. Second, MFTC was annotated by a limited fixed set of annotators for whom we can construct a rich annotation history. However, since there are few annotators per sample to pick from, ACAL cannot leverage this information effectively. Again, we see that strategies perform relatively similarly to one another, except for the scores. Third, in MHS we observe that all strategies using random sample selection require less data. Since the task has low inter-rater agreement scores, uncertainty-based sampling wrongly attempts to sample annotations for correct high-entropy predictions, while this is an accurate distribution.
Our findings highlight that with many labels per sample, ACAL is more data-efficient than traditional AL and passive learning in terms of overall evaluation. However, when data characteristics differ and few annotations are available per sample, ACAL has less of an impact. Polarized agreement scores (either high agreement or no agreement) make the use of ACAL and AL cause little to no improvements over passive learning. This corroborates that (AC)AL leads to improvements in specific cases (Dor et al., 2020). Furthermore, we found conflicting results depending on the metric used ( and ). We closely examine the relationship between the evaluation metrics by turning to annotator-centric evaluation, observing how ACAL impacted predictions for individual annotators.
5.2 Annotator-centric evaluation
We show the annotator-centric evaluation metrics in Tables 2, 3, and 4 for DICES, MFTC (care) and MHS (dehumanize), respectively. We again describe per-dataset results. Again, for DICES and MHS, we observe a positive effect of using ACAL over PL, both in terms of data efficiency and final annotator-centric behavior. For these datasets, ACAL leads to a better representation of annotators on average (,), as well as a better representation of the 10% most different annotators (, ). Compared ACAL to AL, we mainly observe improvements in the DICES dataset, showing less data used and a better annotator-centric score. We observe a strong for the strategy and worse , corroborating our earlier finding that emphasizing diverging label behavior trades off with the averaged evaluation scores. Interestingly, this is not apparent in the scores. For MHS, all approaches using random data sampling () require considerably less data than passive learning. Further, since the pool of annotators for MHS is large (7K+), there will always be some annotators in disagreement with the output of our models, leading to a zero score on .
Average | Worst-off | ||||
App. | |||||
0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
0.442 | 0.186 | 0.164 | 0.447 | -35.5 | |
0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
0.439 | 0.187 | 0.184 | 0.447 | -38.2 | |
0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
0.414 | 0.191 | 0.133 | 0.425 | -0.1 | |
0.384 | 0.192 | 0.117 | 0.427 | -0.1 | |
Passive | 0.371 | 0.211 | 0.123 | 0.479 | – |
Average | Worst-off | ||||
App. | |||||
0.611 | 0.141 | 0.377 | 0.247 | -1.6 | |
0.616 | 0.142 | 0.392 | 0.249 | -0.4 | |
0.600 | 0.145 | 0.351 | 0.248 | -1.7 | |
0.604 | 0.144 | 0.357 | 0.243 | -1.7 | |
0.612 | 0.143 | 0.377 | 0.252 | -5.6 | |
0.589 | 0.142 | 0.423 | 0.248 | -2.5 | |
0.608 | 0.143 | 0.399 | 0.258 | -1.1 | |
0.586 | 0.145 | 0.357 | 0.253 | -2.5 | |
0.586 | 0.141 | 0.392 | 0.255 | -0.2 | |
0.583 | 0.144 | 0.357 | 0.253 | -12.7 | |
Passive | 0.512 | 0.179 | 0.377 | 0.251 | – |
Average | Worst-off | ||||
App. | |||||
0.315 | 0.394 | 0.000 | 0.489 | -50.0 | |
0.322 | 0.397 | 0.000 | 0.478 | -62.5 | |
0.313 | 0.397 | 0.000 | 0.480 | -62.5 | |
0.318 | 0.398 | 0.000 | 0.479 | -62.5 | |
0.322 | 0.389 | 0.000 | 0.508 | -7.8 | |
0.328 | 0.388 | 0.000 | 0.507 | -7.8 | |
0.326 | 0.388 | 0.000 | 0.506 | -7.8 | |
0.326 | 0.384 | 0.000 | 0.513 | -3.0 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
Passive | 0.202 | 0.424 | 0.000 | 0.547 | – |
5.3 Training plots
While the evaluation shows a pattern of efficient data use with ACAL under certain data conditions, it reveals little about how the metrics behave during training or how individual annotator strategies behave. To this end, we provide a complete overview of all metrics (as computed on the validation set) during training in App. B.3. Here we describe the major patterns reoccurring across our experiments using examples and show six of particular interest (Figure 3). Since the strategies only differ in what annotations are included during training, we only show plots related to the annotator-centric metrics.
We can see that there is an influence of both the data sampling and annotator strategies on the performance of the models. Only on DICES is the choice of irrelevant, probably due to the low number of samples. Specifically deteriorates slower for the worst-off annotators than the other strategies but does so without being able to uphold a competitive score. In MFTC, we see that when using , performance on dips at the start of training. Selecting annotators for samples with high predicted entropy initially leads to a decrease in average performance. The strategy seeks to first lower the entropy for the labels already encountered, though some of the variation in labels is irreconcilable. A similar reasoning holds for MHS, where the differences between strategies are even less impacted by the choice of . These two plots further underline our main finding that for ACAL to be impactful in representing diverse annotation perspectives, we need to ensure a (1) heterogeneous pool of annotators, and (2) a task that facilitates human label variation.
5.4 Change in task
In Fig. 4, we present a comparative analysis of two annotator-centric metrics across the three distinct tasks of MFTC and MHS, for which we have seen the least impact of ACAL over AL and PL. We cannot conclude that the chosen ACAL approach () offers a consistent improvement over sampling all annotations (), particularly given that the models using ACAL occasionally require more data to converge (Tables 8 to 11).
Initially, we hypothesized that tasks with a high degree of subjectivity would benefit from ACAL strategies, especially on metrics focused on the most marginalized (worst-off) annotators. These strategies typically involve selecting an annotator whose patterns of annotation diverge from the majority, either in terms of their annotation behavior or in the semantic content of their past annotations.
@@ 5.4 Change in task
However, as depicted in Figure 4, when examining the task of dehumanize (high entropy), it becomes apparent that ACAL does not consistently outperform AL. ACAL demonstrates a lower -score than AL for this task, and on the other hand, a higher -score for a task that is less subjective, such as genocide. Similarly, when evaluating loyalty, which involves the moral dimension with the highest disagreement among annotators, the lower 10% of annotators are more accurately approximated with PL. This suggests that the effectiveness of annotation strategies varies depending on the task’s degree of subjectivity and available pool of annotators. The more heterogeneous the annotation behavior, indicative of a highly subjective task, the larger the pool of annotators required for each item selection. However, due to the limited annotations available per item in both datasets MFTC and MHS, even carefully selecting specific annotators may not adequately represent divergent annotation behavior in general, which challenges the generalization to unseen data. Finally, we can observe that there is a trade-off between modeling the majority of annotators equally, as reflected in the -score and prioritizing the minority viewpoint (). A better performance in one aspect does not necessarily guarantee superiority in the other.
@@ 6 Conclusion
We introduce Annotator-Centric Active Learning (ACAL), an active learning approach that incorporates annotator selection strategies aimed at capturing label variation among annotators. We experiment with tasks across three different datasets, each leading to different ACAL behaviors. One of these datasets, DICES, is the most realistic application of ACAL since the pool of possible annotators is the largest. Here, ACAL leads to more diverse label distributions using fewer annotations. However, we find that the effectiveness of the ACAL paradigm is contingent on data characteristics. These characteristics include the number of annotations per sample, the number of annotations per annotator, and the nature of disagreement in the task annotations. Our analysis shows that we can use these conditions to help explain the often disappointing results for AL in NLP applications.
Including annotator-centric evaluation reveals how methods with similar averaged performance deal with different levels of disagreement among annotators. We show that evaluation can be enhanced by focusing on individual annotators, as there is a large gap between conventional, averaged, and worst-off performance. Furthermore, many aspects of our ACAL approach can be experimented with, e.g. by swapping the order in which samples are selected (in our case first) and annotators (second), or investigating the impact of including annotator-specific demographic information, as it is inconsistently predictive of annotation behavior Orlikowski et al. (2023); Beck et al. (2024).
@@
The main limitation of this work is that the experiments are based on simulated active learning which is known to bear potential issues (Margatina and Aletras, 2023). In our study, a primary challenge arises with two of the datasets (MFTC, MHS), which, despite having a large pool of annotators, lack annotations from every annotator for each item. Consequently, in real-world scenarios, the annotator selection strategies for these datasets would benefit from access to a more extensive pool of annotators. This limitation likely contributes to the underperformance of ACAL on these datasets compared to DICES. We emphasize the need for more datasets that feature a greater number of annotations per item, as this would significantly enhance research efforts aimed at modeling human disagreement.
Since we evaluate four different annotator selection strategies and two sample selection strategies across three datasets and seven tasks, the amount of experiments is high. This did not allow for further investigation of the difference using different classification models, the extensive turning of hyperparameters, or even different training paradigms. Lastly, a limitation of our annotator selection strategies is that they rely on a small annotation history. This is why we require a warmup phase for some of the strategies, for which we decided to take a random sample of annotations. Incorporating more informed warmup strategies or incorporating ACAL strategies that do not rely on annotator history may positively impact its performance and data efficiency.
@@ Appendix A Detailed Experimental Setup
A.1 Cross validation details
We split the data on samples, meaning that all annotations for any given sample are completely contained in each separate split.
@@ A.2 Hyperparameters
We report the hyperparameters for training passive, AL, and ACAL in Tables 5, 6, and 7, respectively. For turning the learning rate for passive learning, on each dataset, we started with a learning rate of 1e-06 and increased it by a factor of 3 in steps until the model showed a tendency to overfit quickly (within a single epoch). All other parameters are kept on their default setting.
@@
num rounds | DICES | 50 |
num rounds | MFTC (all), MHS (all) | 20 |

### v2

@@ 5 Results
We start by highlighting the benefits of ACAL over AL and PL (Section 5.1). Next, we closely examine ACAL on efficiency and fairness (Section 5.2). Then, we select a few cases of interest and dive deeper into the strategies’ behavior during training (Section 5.3). Finally, we investigate ACAL across varying levels of subjectivity (Section 5.4).
[width=.5]tikz/snapshot
5.1 Highlights
Our experiments show that ACAL can have a beneficial impact over using PL and AL. Figure 2 highlights two main findings: (1) ACAL strategies can more quickly learn to represent the annotation distribution with a large pool of annotators, and (2) when agreement between annotators is polarized, ACAL leads to improved results compared to learning from aggregated labels. In the next sections, we provide a deeper understanding of the conditions in which ACAL works well.
5.2 Efficiency and Fairness
Table 1 presents the results of evaluating the best models on the test set. We analyze the results along two dimensions: (a) efficiency: what is the impact of the different strategies on the trade-off between annotation budget and performance? (b) fairness: do the selection strategies that aim for a balanced consideration of minority and majority views lead to better performance in the human-centric evaluation metrics? For MFTC we focus on care because it has an average number of samples available, and for MHS we focus on dehumanize because it has high levels of disagreement. Appendix B presents additional results.
Average | Worst-off | |||||||
App. | ||||||||
DICES | 53.2 | .100 | 43.2 | .186 | 16.7 | .453 | -36.8 | |
55.5 | .101 | 42.4 | .187 | 15.5 | .450 | -32.7 | ||
61.0 | .103 | 44.2 | .186 | 16.4 | .447 | -35.5 | ||
58.9 | .142 | 43.1 | .203 | 16.9 | .370 | -30.0 | ||
53.2 | .100 | 43.2 | .186 | 16.7 | .453 | -36.8 | ||
55.5 | .101 | 42.4 | .187 | 15.5 | .450 | -32.7 | ||
63.1 | .098 | 43.9 | .187 | 18.4 | .447 | -38.2 | ||
58.9 | .142 | 43.1 | .203 | 16.9 | .370 | -30.0 | ||
59.1 | .112 | 41.4 | .191 | 13.3 | .425 | -0.1 | ||
46.2 | .110 | 38.4 | .192 | 11.7 | .427 | -0.1 | ||
PL | 59.0 | .105 | 37.1 | .211 | 12.3 | .479 | – | |
MFTC (care) | 78.9 | .038 | 61.1 | .141 | 37.7 | .247 | -1.6 | |
78.5 | .037 | 61.6 | .142 | 39.2 | .249 | -0.4 | ||
78.1 | .039 | 60.0 | .145 | 35.1 | .248 | -1.7 | ||
76.6 | .040 | 60.4 | .144 | 35.7 | .243 | -1.7 | ||
79.4 | .038 | 61.2 | .143 | 37.7 | .252 | -5.6 | ||
80.7 | .037 | 58.9 | .142 | 42.3 | .248 | -2.5 | ||
79.1 | .037 | 60.8 | .143 | 39.9 | .258 | -1.1 | ||
78.1 | .040 | 58.6 | .145 | 35.7 | .253 | -2.5 | ||
79.0 | .037 | 58.6 | .141 | 39.2 | .255 | -0.2 | ||
79.4 | .037 | 58.3 | .144 | 35.7 | .253 | -12.7 | ||
PL | 81.1 | .032 | 51.2 | .179 | 37.7 | .251 | – | |
MHS (dehumanize) | 33.6 | .081 | 31.5 | .394 | 0.0 | .489 | -50.0 | |
33.1 | .081 | 32.2 | .397 | 0.0 | .478 | -62.5 | ||
30.5 | .079 | 31.3 | .397 | 0.0 | .480 | -62.5 | ||
32.4 | .081 | 31.8 | .398 | 0.0 | .479 | -62.5 | ||
32.4 | .080 | 32.2 | .389 | 0.0 | .508 | -7.8 | ||
33.1 | .080 | 32.8 | .388 | 0.0 | .507 | -7.8 | ||
33.6 | .080 | 32.6 | .388 | 0.0 | .506 | -7.8 | ||
33.0 | .079 | 32.6 | .384 | 0.0 | .513 | -3.0 | ||
32.8 | .077 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
PL | 28.0 | .075 | 20.2 | .424 | 0.0 | .547 | – |
Efficiency
We discuss the performance on and to measure how well the proposed strategies model label distributions and examine the used annotator budget. Across all tasks and datasets, ACAL and AL consistently yield comparable or superior and with a lower annotation budget than PL. When comparing ACAL with AL, the results vary depending on the task and dataset. For DICES, there is a significant benefit to using ACAL, as it can save up to 40% of the annotation budget while yielding better scores across all metrics than AL. With AL, we observe only a small reduction in annotation cost. For MFTC, AL with leads to the largest cost benefits (12% less annotation budget), but at a cost in terms of absolute and . ACAL slightly outperforms AL but does not lead to a decrease in annotation budget. For MHS, both AL and ACAL significantly reduce the annotation cost (60%) while yielding better scores than PL—however, AL and ACAL do not show substantial performance differences.
Overall, we conclude that ACAL is most efficient when the pool of available annotators for one sample is large (as with the DICES dataset), whereas the difference between ACAL and AL is negligible with a small pool of annotators per data sample (as with MFTC and MHS).
Fairness
We investigate the extent to which the models represent individual annotators fairly and capture minority opinions via the annotator-centric evaluation metrics (, , , and ). We observe a substantial improvement when using AL or ACAL over PL. Further, we observe no single winner-takes-all approach: high and scores do not consistently cooccur with high scores for the annotator-centric metrics. This highlights the need for a more comprehensive evaluation to assess models for subjective tasks. We observe that ACAL slightly outperforms AL in modeling individual annotators ( and ). This trend is particularly evident with DICES, again likely due to the large pool of annotators available per data sample. Lastly, ACAL is best in the worst-off metrics ( and ), showing the ability to better represent minority opinions as a direct consequence of the proposed annotator selection strategies on DICES and MFTC. However, all approaches score 0 for on MHS. This is due to the high disagreement in this dataset: the 10% worst-off annotators always disagree with a hard label derived from the predicted label distribution.
In conclusion, our experiments show that, when a large pool of annotators is available, a targeted sampling of annotators requires fewer annotations and is fairer. That is, minority opinions are better represented without large sacrifices in performance compared to the overall label distribution.
[width=]tikz/training_process_selected
5.3 Convergence
The evaluation on the test set paints a general picture of the advantage of using ACAL over AL or PL. In this section, we assess how different ACAL strategies converge over iterations. We describe the major patterns across our experiments by analyzing six examples of interest with and (Figure 3). We select because it reveals how well individual annotators are modeled on average, and to measure how strategies deviate from modeling the majority perspective. Appendix B.2 provides an overview of all metrics.
First, we notice that the trends for and are both increasing—the first is expected, but the second requires an explanation. As the model is exposed to more annotations over the training iterations, the predicted label distribution starts to fit the true label distribution. However, here we consider each annotator individually: reports the average of the 10% lowest scores per annotator. The presence of disagreement implies the existence of annotators that annotate differently from the majority. Since our models predict the full distribution, they assign a proportional probability to dissenting annotators. Thus, learning to model the full distribution of annotations leads to an increase in .
Second, we notice a difference between ACAL and AL. On MFTC and MHS, ACAL, compared to AL, yields overall smaller at the cost of a slower convergence in , showing the trade-off between modeling all annotators and representing minorities. However, with DICES the trend is the opposite. This is due to AL having access to the complete label distribution: it can model a balanced distribution, leading to lower worst-off performance. With a large number of annotations, ACAL requires more iterations to get the same balanced predicted distribution.
Third, we observe differences among the annotator selection strategies (). shows the most differences—both and increase slower than for the other strategies. This suggests that selecting annotators based on the average embedding of the annotated content strongest emphasizes diverging label behavior.
Finally, we analyze the impact of the sample selection strategies (, dotted vs. solid lines in Figure 3). For DICES, and lead to comparable results, likely due to the low number of samples. Using in MFTC leads to performance decreasing at the start of training. The strategy prioritizes obtaining annotations for already added samples to lower their entropy, while the variation in labels is irreconcilable (since there are limited labels available, and they are in disagreement). We see a similar pattern for MHS.
These results further underline our main finding that ACAL is effective in representing diverse annotation perspectives when there is a (1) heterogeneous pool of annotators, and (2) a task that facilitates human label variation.
5.4 Impact of subjectivity
We further investigate ACAL strategies on (1) label entropy, and (2) cross-task performance.
Alignment of ACAL strategies during training
We want to investigate how well the ACAL strategies align with the overall subjective annotations: do they drive the model entropy in the right direction? We measure the entropy of the samples in the labeled training set at each iteration and compare it to the actual entropy of those samples. Higher entropy suggests that the selection strategy overestimates uncertainty. Lower entropy indicates that the model may not sufficiently account for disagreement. When the entropy matches the true entropy, the selection strategy is well-calibrated. We focus on DICES as a case study due to the wide range of entropy scores. We group each sample based on the true label entropy into low, medium, and high111Entropy bins: low (), medium () high ().. We apply the same categorization at each training iteration for samples labeled thus far. Subsequently, we plot the proportion of data points for which the selection strategy results in excessively high or excessively low entropy.
[width=]tikz/align_training
Figure 4 visualizes the proportions. At the beginning of training, entropy is generally low because samples have few annotations. Over time, the selected annotations better align with the true entropy. However, when and how much strategies succeeded in representing the true label distribution differs: and take longer to increase label entropy than the other two strategies. They are conservative in adding diverse labels. and increase the proportion of well-aligned data points earlier in the training process, achieving a balanced entropy alignment sooner. However, both strategies start to overshoot the target entropy, whereas the others show a more gradual alignment with the true entropy. This effect is strongest for . This finding suggests that minority-aware annotator-selection strategies achieve the best results in the early stages of training. They are effective for quickly raising entropy but can lead to overrepresentation.
Cross-task performance
Figure 5 compares the two annotator-centric metrics on the three tasks of MFTC and MHS—the datasets for which we have seen the least impact of ACAL over AL and PL. We select a data sampling () and annotator sampling strategy (), based on its strong performance on DICES for comprehensive comparison.
[width=]tikz/comparison_tasks_SRTS
@@ 5.4 Impact of subjectivity
When evaluating MFTC loyalty, which has the highest disagreement, is more accurately approximated with PL. Similarly, ACAL is outperformed by AL on for the dehumanize (high disagreement) task. However, for the less subjective task genocide, ACAL leads to higher . This suggests that the effectiveness of annotation strategies varies depending on the task’s degree of subjectivity and the available pool of annotators. The more heterogeneous the annotation behavior, indicative of a highly subjective task, the larger the pool of annotators required for each sample selection. We also observe that there is a trade-off between modeling the majority of annotators equally () and prioritizing the minority ().
@@ 6 Conclusion
We present ACAL as an extension of AL to emphasize the selection of diverse annotators. We introduce three novel annotator selection strategies and four annotator-centric metrics and experiment with tasks across three different datasets. We find that the ACAL approach is especially effective in reducing the annotation budget when the pool of available annotators is large. However, its effectiveness is contingent on data characteristics such as the number of annotations per sample, the number of annotations per annotator, and the nature of disagreement in the task annotations. Furthermore, our novel evaluation metrics display the trade-off between modeling overall distributions of annotations and adequately accounting for minority voices, showing that different strategies can be tailored to meet different goals. Especially early in the training process, strategies that are aggressive in obtaining diverse labels have a beneficial impact Furthermore, we recognize that gathering a distribution that uniformly contains all possible (minority and majority) labels can be overly sensitive to small minorities or noise. Future work can integrate methods that account for noisy annotations Weber-Genzel et al. (2024) or that strike a balance between egalitarian and utilitarian approaches Lera-Leri et al. (2024).
@@
The main limitation of this work is that the experiments are based on simulated AL which is known to bear potential issues (Margatina and Aletras, 2023). In our study, a primary challenge arises with two of the datasets (MFTC, MHS), which, despite having a large pool of annotators, lack annotations from every annotator for each item. Consequently, in real-world scenarios, the annotator selection strategies for these datasets would benefit from access to a more extensive pool of annotators. This limitation likely contributes to the underperformance of ACAL on these datasets compared to DICES. We emphasize the need for more datasets that feature a greater number of annotations per item, as this would significantly enhance research efforts aimed at modeling human disagreement.
Since we evaluate four different annotator selection strategies and two sample selection strategies across three datasets and seven tasks, the amount of experiments is high. This did not allow for further investigation of other methods for measuring uncertainty (such as ensemble methods ), different classification models, the extensive turning of hyperparameters, or even different training paradigms (such as low-rank adaptation ). Lastly, a limitation of our annotator selection strategies is that they rely on a small annotation history. This is why we require a warmup phase for some of the strategies, for which we decided to take a random sample of annotations. Incorporating more informed warmup strategies or incorporating ACAL strategies that do not rely on annotator history may positively impact its performance and data efficiency.
@@ Appendix A Detailed Experimental Setup
Dataset | Task (dimension) | # Samples | # Annotators | # Annotations | # Annotations per item |
DICES | Safety Judgment | 990 | 172 | 72,103 | 72.83 |
MFTC | Morality (care) | 8,434 | 23 | 31,310 | 3.71 |
MFTC | Morality (loyalty) | 3,288 | 23 | 12,803 | 3.89 |
MFTC | Morality (betrayal) | 12,546 | 23 | 47,002 | 3.75 |
MHS | Hate Speech (dehumanize, genocide, respect) | 17,282 | 7,807 | 57,980 | 3.35 |
A.1 Dataset details
We provide an overview of the datasets used in our work in Table A1. We split the data on samples, meaning that all annotations for any given sample are completely contained in each separate split.
@@ A.2 Hyperparameters
We report the hyperparameters for training passive, AL, and ACAL in Tables A2, A3, and A4, respectively. For turning the learning rate for passive learning, on each dataset, we started with a learning rate of 1e-06 and increased it by a factor of 3 in steps until the model showed a tendency to overfit quickly (within a single epoch). All other parameters are kept on their default setting.
@@
num iterations | DICES | 50 |
num iterations | MFTC (all), MHS (all) | 20 |